# HalluVault: A Novel Logic Programming-aided Metamorphic Testing Framework for Detecting Fact-Conflicting Hallucinations in Large Language Models

### v1

@@ Abstract.
Large language models (LLMs) have transformed the landscape of language processing, yet struggle with significant challenges in terms of security, privacy, and the generation of seemingly coherent but factually inaccurate outputs, commonly referred to as hallucinations. Among these challenges, one particularly pressing issue is Fact-Conflicting Hallucination (FCH), where LLMs generate content that directly contradicts established facts. Tackling FCH poses a formidable task due to two primary obstacles: Firstly, the hard task of automatically constructing and updating benchmark datasets, since current methods heavily rely on manually curated static benchmarks that fail to encompass the wide-ranging and evolving spectrum of FCH scenarios. Secondly, validating the reasoning process behind LLM outputs is inherently complex, especially when it involves intricate logical relations.
In addressing these obstacles, we propose an innovative approach leveraging logic programming to enhance metamorphic testing for detecting Fact-Conflicting Hallucinations (FCH). Our method introduces a sophisticated framework that leverages crawling techniques to gather information from sources such as Wikipedia, seamlessly integrated within our tool. By employing logical reasoning principles, we refine and expand this knowledge base to generate a diverse array of test cases with definitive answers. These test cases are then employed to assess LLMs through structured prompts, prompting them to deliver responses backed by logical reasoning. To validate the coherence of their reasoning, we present two semantic-aware assessment mechanisms that evaluate the alignment between the logical and semantic structures of the LLM responses and the ground truth. Our method autonomously generates pertinent test cases and detects hallucinations across six different LLMs spanning nine domains, revealing hallucination rates ranging from 24.7% to 59.8%. Key observations indicate that LLMs encounter challenges particularly with temporal concepts, handling out-of-distribution knowledge, and exhibiting deficiencies in logical reasoning capabilities. The outcomes underscore the efficacy of logic-based test cases generated by our tool in both triggering and identifying hallucinations. To further mitigate the identified FCHs, we delve into model editing techniques, which demonstrated effectiveness on a limited scale (involving edits to fewer than 1000 knowledge pieces). These findings underscore the imperative for ongoing collaborative endeavors within the community to detect and address LLM hallucinations.
@@ 1. Introduction
Large Language Models (LLMs) have revolutionized the landscape of language processing and beyond, exhibiting remarkable capabilities in text generation and comprehension with a wide range of applications. However, despite their growing prominence, LLMs encounter significant challenges in terms of security and privacy, which greatly impact their effectiveness and reliability. One particularly notable issue is the phenomenon of “hallucination”, where LLMs generate coherent but factually incorrect or irrelevant outputs during tasks such as problem-solving. This tendency to produce misleading information not only compromises the safety of LLM applications but also raises serious concerns about their usability. Among the various forms of hallucinations in LLMs, “Fact-conflicting hallucination” (FCH) is a primary concern and the main focus of this paper. FCH occurs when LLMs generate content that directly contradicts established facts, as illustrated in Figure 2. For example, an LLM might incorrectly believe that “Haruki Murakami won the Nobel Prize in Literature in 2016”, deviating from the correct answer that “Haruki Murakami has not won the Nobel Prize but has received numerous other awards for his work in Literature”. Such misinformation dissemination leads to significant user confusion, eroding the trust and reliability that are essential in various LLM applications.
@@
To tackle the issue of hallucinations in LLMs, recent studies have introduced various methods for their detection and testing. A common and straightforward approach involves creating extensive benchmarks tailored for this purpose. Datasets such as TruthfulQA, HaluEval, and KoLA have been designed to evaluate hallucinations across different contexts, including question-answering, summarization, and knowledge graphs. However, the current research landscape in LLM presents a critical gap in automatically and effectively testing FCHs. The main obstacle in testing for FCH is the absence of dedicated ground truth datasets and specific testing frameworks. Unlike other types of hallucinations (e.g., input-conflicting and context-conflicting hallucinations, to be detailed in Section 2), which can be identified through checks for semantic consistency, FCH requires the verification of the content’s factual accuracy against external sources of knowledge or databases. This requirement makes the process particularly challenging and resource-intensive, especially for tasks processing contents with inherent logical connections.
@@
Bridging the identified research gap in the literature necessitates an exploration of the inherent challenges faced in detecting FCHs, which are crucial for advancing and enhancing the reliability of LLMs. The first challenge is the difficulty in automatically constructing and updating benchmark datasets. Predominantly, existing methodologies rely on manually curated benchmarks. While these benchmarks are effective in detecting certain types of hallucinations, they fall short in encompassing the broad and dynamic spectrum of fact-conflicting scenarios inherent to LLMs. Moreover, the need for frequent updates to benchmark data, due to the ever-evolving nature of knowledge, imposes a significant and continuous maintenance effort. The reliance on benchmark datasets thus restricts the detection techniques’ adaptability, scalability, and worse, detection capability. The second challenge is the difficulty in automatically validating answers from LLM outputs. Even when LLMs produce correct final answers, the outputs may not represent the true reasoning process behind them, potentially masking false understanding – a source of FCH hallucination. Automatically validating the reasoning process, especially those involving complex logic relations, is inherently difficult. Furthermore, the consistency in the quality of benchmark questions can vary due to the differing levels of experience and skill among human experts creating them, introducing noise, particularly in data labeling and result validation stages.
@@
To address limitations in the existing techniques, we are the first, to the best of our knowledge, to introduce a novel automatic logic-programming-aided metamorphic testing technique for hallucination detection in this work. We have developed an extensive and extensible FCH testing framework, which is based on factual knowledge reasoning and metamorphic testing, seamlessly integrated into our tool called HalluVault. HalluVault begins by establishing a comprehensive factual knowledge base, sourced through extensive crawling of information from accessible knowledge bases such as Wikipedia. Each piece of this knowledge acts as a “seed” for subsequent transformations. Leveraging logic reasoning relations, we transform and augment these seeds, thereby expanding the factual knowledge into a well-established set of question-answer pairs. Using the questions and answers in the knowledge set as test cases and ground truth respectively, we construct a reliable and robust FCH testing benchmark. This is implemented through a series of well-formulated template-based prompts to test FCH in LLMs. Specifically, we instruct the LLMs to generate their answers to the test cases. To facilitate a thorough evaluation of the reasoning logic behind their responses, we require the LLMs to provide detailed justifications for their answers. For effective and dependable identification of FCH, we introduce two semantic-aware and similarity-based metamorphic oracles. These oracles operate by extracting essential semantic elements from each sentence and mapping out their logical relationships. By assessing the similarity between the constructed logical and semantic structures of the LLM’s answers and the ground truth, we can detect FCH by pinpointing answers that significantly diverge from the ground truth.
@@
In evaluating our proposed FCH testing framework and HalluVault, we undertake comprehensive experiments to evaluate their effectiveness in a wide array of contexts. On the one hand, our evaluation strategy involves deploying HalluVault across a broad spectrum of topics, sourced from an extensive and diverse range of Wikipedia articles. On the other hand, we examine our framework on a variety of open-source and commercial LLMs, providing a thorough examination of its applicability and performance across different model architectures.
@@
Our key findings indicate that HalluVault succeeds in automatically generating useful test cases and identifying hallucination issues of six LLMs across nine domains. Using these test sets, we find that hallucination responses generated by different LLMs can vary from 24.7% to 59.8%. We then categorize these hallucination responses into four types. Through an in-depth analysis, we unveil that the lack of logical reasoning capabilities contributes the most to the FCH issues in LLMs. Additionally, we observe that LLMs are particularly prone to generating hallucinations in test cases involving temporal concepts and out-of-distribution knowledge. Furthermore, we confirm that test cases generated using our logical reasoning rules can effectively trigger and detect hallucination issues in LLMs. As mitigation, we investigate the use of model editing techniques to rectify the identified FCHs. These techniques have shown promising results when applied on a small scale (involving edits up to less than 1000 pieces of knowledge). Our results highlight the importance of ongoing efforts within the community to detect and address issues of hallucination in LLMs.
The main contributions of this paper are as follows:
@@
Development of a novel FCH Testing Framework: To the best of our knowledge, we are the first to develop a novel testing framework based on logic programming and metamorphic testing to automatically detect FCH issues in LLMs. This framework represents a significant advancement over current methodologies, providing a more systematic, comprehensive approach to detection.
@@
Construction and Release of Extensive Factual Knowledge Base and Benchmark: Our work constructs a large-scale benchmark dataset to facilitate collaborative efforts and future advancements in the detection of FCH.
@@
Designing and Implementing Innovative Logic-reasoning-based Method for Data Mutation: We propose and implement five unique logic reasoning rules to mutate and augment the initial seeds from our knowledge base, increasing the diversity and effectiveness of our test scenario.
@@
Deployment of FCH-specific semantic-aware testing oracles for automatic LLM answer validation: We propose and implement two automated verification mechanisms (oracles) that analyze the semantic structure similarity between sentences. These oracles are designed to validate the reasoning logic behind the answers generated by LLMs, hereby reliably detecting the occurrence of FCHs.
@@ 2.1. Hallucination Categorization
Hallucination in LLMs can be categorized into the three main categories (Yao et al., 2023; Huang et al., 2023; Zhang et al., 2023), as detailed below.
@@ 2.2. Logic Programming
Logic programming languages are declarative, i.e., programming with these languages means describing the world. Using the programs means asking questions about the previously described world. Based on the answers to the questions from the logic reasoning engine, according to the world description, we can acquire new facts. Figure 2 shows an example of how logic programming works. Here we explain some key concepts:
@@
An example is , which means horses are bigger than cats. Another example is , which means that Gunzon Prize is in the list of prizes awarded to Haruki Murakami.
Rules. A rule is a statement defining the logical relations of two or more facts and rules. It is made up of a (a new predicate) and a (a sequence of facts or rules separated by commas). It is denoted as:
An example is , which means the relation is transitive. Another example is , which means is an inverse relation of . The last example here is , which is if there exists at least one that is a member of both and . It is a composition of two predicates.
Queries. A query has the same structure as the body of a rule, i.e., it is a sequence of predicates separated by commas. It is denoted as:
The logic reasoning engine will answer if the sequence of predicates in the query is according to the facts and rules. Otherwise, it will answer . An example query is , which means asking the logic reasoning engine whether elephants are bigger than cats. Another example is , which means asking if the awards won by Haruki Murakami and Bob Dylan have overlaps.
Reasoning Rules. As shown in Figure 2, generating new facts through logic programming requires facts (Equation (missing) 1), rules (Equation (missing) 2), queries (Equation (missing) 3), and answers to the queries. To simplify the notation of this process, we bring up the concept of reasoning rules in this paper, which describes the inference process of using facts and rules (predicates) to reach the conclusion (a new fact in the form of a predicate) by omitting the process of querying and analyzing the query answers. A reasoning rule is denoted in this form:
@@ 3. Motivating Example
Figure 3 shows a motivating example of HalluVault. Assume we have the facts about whether Haruki Murakami and Bob Dylan have won the Nobel Prize, as illustrated in the left sub-figure. The question to ask LLMs is straightforward: We can ask whether Haruki Murakami/Bob Dylan has won the Nobel Prize or not. Asking and verifying this knowledge require no logic reasoning. However, the straightforward questions are often not enough to unveil hallucinations. Therefore, more diversified questions (questions with intertwined and complex information, as illustrated in the right sub-figure) are needed.
In order to generate more diversified benchmarks, previous research (Yu et al., 2023; Li et al., 2023a) involves human experts to generate the questions and annotate the answers for hallucination checking. Although the manually generated benchmarks can unveil certain hallucinations, they suffer from several drawbacks. The landscape of knowledge is dynamic, with new information continuously surfacing and older information becoming obsolete. If facts change continuous over time, for instance Haruki Murakami were to win the Nobel Prize in the future, this would necessitate regular updates and corrections to the ground truth in existing datasets to reflect them. However, maintaining the accuracy of these benchmarks demands a significant amount of manual labor. Additionally, the quality of the questions in the benchmarks might be inconsistent due to the differences in the experience and skills of the human experts who create them. Consequently, the efficiency and soundness of the manually generated benchmarks does not have a guarantee.
@@
These two challenges can both be addressed by leveraging logic programming. We can derive new logically sound facts based on existing knowledge. With the new facts, we can then generate diverse questions and their ground truth answers. With the ground truth answers, we can generate test oracles to capture hallucinations. In short, the idea of using logic programming to tackle the challenges motivates the design of HalluVault.

### v2

# Drowzee: Metamorphic Testing for Fact-Conflicting Hallucination Detection in Large Language Models

@@ Abstract.
Large language models (LLMs) have revolutionized language processing, but face critical challenges with security, privacy, and generating hallucinations — coherent but factually inaccurate outputs. A major issue is fact-conflicting hallucination (FCH), where LLMs produce content contradicting ground truth facts. Addressing FCH is difficult due to two key challenges: 1) Automatically constructing and updating benchmark datasets is hard, as existing methods rely on manually curated static benchmarks that cannot cover the broad, evolving spectrum of FCH cases. 2) Validating the reasoning behind LLM outputs is inherently difficult, especially for complex logical relations.
To tackle these challenges, we introduce a novel logic-programming-aided metamorphic testing technique for FCH detection. We develop an extensive and extensible framework that constructs a comprehensive factual knowledge base by crawling sources like Wikipedia, seamlessly integrated into Drowzee111Drowzee is named after a Pokémon (Satoshi Tajiri, 2023) character that nourishes itself by eating dreams. This name symbolizes our tool’s capability to detect and potentially further assist in eliminating the hallucinations in LLMs.. Using logical reasoning rules, we transform and augment this knowledge into a large set of test cases with ground truth answers. We test LLMs on these cases through template-based prompts, requiring them to provide reasoned answers. To validate their reasoning, we propose two semantic-aware oracles that assess the similarity between the semantic structures of the LLM answers and ground truth. Our approach automatically generates useful test cases and identifies hallucinations across six LLMs within nine domains, with hallucination rates ranging from 24.7% to 59.8%. Key findings include LLMs struggling with temporal concepts, out-of-distribution knowledge, and lack of logical reasoning capabilities. The results show that logic-based test cases generated by Drowzee effectively trigger and detect hallucinations. To further mitigate the identified FCHs, we explored model editing techniques, which proved effective on a small scale (with edits to fewer than 1000 knowledge pieces). Our findings emphasize the need for continued community efforts to detect and mitigate model hallucinations.
@@ 1. Introduction
Large Language Models (LLMs) have brought transformative advancements to the fields of language processing and beyond, showcasing exceptional abilities in text generation and comprehension with wide-ranging applications. However, despite their increasing prevalence, LLMs face critical challenges in security and privacy aspects (Siddiq and Santos, 2023; Hou et al., 2023; Kaddour et al., 2023; Yang et al., 2024a; Zhang et al., 2024; Xu et al., 2024), heavily impacting their effectiveness and reliability. A particularly notable issue among these is the phenomenon of “hallucination”, where LLMs produce coherent but factually inaccurate or irrelevant outputs during tasks like problem-solving. This tendency to generate misleading information not only jeopardizes the safety of LLM applications but also raises serious usability concerns. Hallucinations in LLMs take several forms, with “Fact-conflicting hallucination” (FCH) being a major concern and the primary type of concern in this paper. FCH is manifested by LLMs generating content that directly contradicts established facts, as exemplified in Figure 1. When an LLM incorrectly believes “Haruki Murakami won the Nobel Prize in Literature in 2016”, deviating from the correct answer of “Haruki Murakami has not won the Nobel Prize but other numerous awards for his work in Literature”. Such misinformation dissemination leads to significant user confusion, eroding the trust and reliability that are crucial in various LLM applications.
@@
To address the issue of hallucinations in LLMs, recent studies have introduced a range of methods for their detection and testing. A common and straightforward approach involves creating extensive benchmarks tailored for this purpose. Datasets such as TruthfulQA (Lin et al., 2022), HaluEval (Li et al., 2023a), and KoLA (Yu et al., 2024) have been designed to evaluate hallucinations across different contexts, including question-answering, summarization, and knowledge graphs. Despite the value of these manually labeled datasets, the current techniques for hallucination detection and testing heavily rely on naive and semi-automatic methods, such as string matching, manual validation, or utilizing another LLM for confirmation. This current research landscape in LLM, however, presents a critical gap in automatically and effectively testing FCHs. The main obstacle in testing for FCH is the absence of dedicated ground truth datasets and specific testing frameworks. Unlike other types of hallucinations (e.g., input-conflicting and context-conflicting hallucinations, to be detailed in Section 2.1) which can be identified through checks for semantic consistency, FCH demands the verification of the content’s factual accuracy against external sources of knowledge or databases. This requirement makes the process particularly challenging and resource-intensive, especially for tasks processing contents with inherent logical connections.
@@
Bridging the identified research gap in the literature necessitates an exploration of the inherent challenges faced in detecting FCHs, which are crucial for advancing and enhancing the reliability of LLMs. Challenge#1: difficulty in automatically constructing and updating benchmark datasets. Predominantly, existing methodologies are anchored to manually curated benchmarks. While these benchmarks are effective in detecting certain types of hallucinations, they fall short in encompassing the broad and dynamic spectrum of fact-conflicting scenarios inherent to LLMs. Meanwhile, the need for frequent updates to benchmark data, due to the ever-evolving nature of knowledge, imposes a significant and continuous maintenance effort. The reliance on benchmark datasets thus restricts the detection techniques’ adaptability, scalability, and worse, detection capability. Challenge#2: difficulty in automatically validating answers from LLM outputs. Even when LLMs produce correct final answers, the outputs may not represent the true reasoning process behind them, potentially masking false understanding – a source of FCH hallucination. Automatically validating the reasoning process, especially those involving complex logic relations, is inherently difficult. Furthermore, the consistency in the quality of benchmark questions can vary due to the differing levels of experience and skill among human experts creating them, introducing noise, particularly in data labeling and result validation stages.
Our Work. To address limitations in the existing techniques, we are the first, to the best of our knowledge, to introduce a novel automatic logic-programming-aided metamorphic testing technique for hallucination detection in this work. We have developed an extensive and extensible FCH testing framework, which is based on factual knowledge reasoning and metamorphic testing, seamlessly integrated into Drowzee.
@@
Drowzee begins by establishing a comprehensive factual knowledge base, sourced through extensive crawling of information from accessible knowledge bases such as Wikipedia. Each piece of this knowledge acts as a “seed” for subsequent transformations. Leveraging logic reasoning relations, we transform and augment these seeds, thereby expanding the factual knowledge into a well-established set of question-answer pairs. Using the questions and answers in the knowledge set as test cases and ground truth respectively, we construct a reliable and robust FCH testing benchmark. This is implemented through a series of well-formulated template-based prompts to test FCH in LLMs. Specifically, we instruct the LLMs to generate their answers to the test cases. To facilitate a thorough evaluation of the reasoning logic behind their responses, we require the LLMs to provide detailed justifications for their answers. For effective and dependable identification of FCH, we introduce two semantic-aware and similarity-based metamorphic oracles. These oracles operate by extracting essential semantic elements from each sentence and mapping out their logical relationships. By assessing the similarity between the constructed logical and semantic structures of the LLM’s answers and the ground truth, we can detect FCH by pinpointing answers that significantly diverge from the ground truth.
@@
Results and Findings. In evaluating our proposed FCH testing framework and Drowzee, we undertake comprehensive experiments to evaluate their effectiveness in a wide array of contexts. On the one hand, our evaluation strategy involves deploying Drowzee across a broad spectrum of topics, sourced from an extensive and diverse range of Wikipedia articles. On the other hand, we examine our framework on a variety of open-source and commercial LLMs, providing a thorough examination of its applicability and performance across different model architectures.
@@
Our key findings indicate that Drowzee succeeds in automatically generating useful test cases and identifying hallucination issues of six LLMs across nine domains. Using these test sets, we find that hallucination responses generated by different LLMs can vary from 24.7% to 59.8%. We then categorize these hallucination responses into four types. Through an in-depth analysis, we unveil that the lack of logical reasoning capabilities contributes the most to the FCH issues in LLMs. Additionally, we observe that LLMs are particularly prone to generating hallucinations in test cases involving temporal concepts and out-of-distribution knowledge. Furthermore, we confirm that test cases generated using our logical reasoning rules can effectively trigger and detect hallucination issues in LLMs. As mitigation, we investigate the use of model editing techniques (Meng et al., 2022; hiyouga, 2023) to rectify the identified FCHs. These techniques have shown promising results when applied on a small scale (involving edits up to less than 1000 pieces of knowledge). Our results highlight the importance of ongoing efforts within the community to detect and address issues of hallucination in LLMs.
Contributions. We summarize the main contributions of this paper below:
@@
Development of a novel FCH Testing Framework. To the best of our knowledge, we are the first to develop a novel testing framework based on logic programming and metamorphic testing to automatically detect FCH issues in LLMs.
@@
Construction and Release of Extensive Factual Knowledge Base and Benchmark. Our work constructs a large-scale benchmark dataset (GitHub, 2024) to facilitate collaborative efforts and future advancements in the detection of FCH.
@@
Designing and Implementing Innovative Logic-reasoning-based Method for Data Mutation. We propose and implement five unique logic reasoning rules to mutate and augment the initial seeds from our knowledge base, increasing the diversity and effectiveness of our test scenario.
@@
Deployment of FCH-specific semantic-aware testing oracles for automatic LLM answer validation. We propose and implement two automated verification mechanisms (oracles) that analyze the semantic structure similarity between sentences. These oracles are designed to validate the reasoning logic behind the answers generated by LLMs, hereby reliably detecting the occurrence of FCHs.
@@ 2.1. Hallucination Categorization
Hallucination in LLMs can be categorized into three main categories (Yao et al., 2024; Huang et al., 2023; Zhang et al., 2023), as detailed below.
@@ 2.2. Logic Programming
Some existing works (Ye et al., 2023; Pan et al., 2023; Olausson et al., 2023) have already integrated logical programming with large language models in an attempt to enhance their logical reasoning capabilities. In this work, we focus on leveraging logical programming to automate the testing of hallucinations in LLMs. Logic programming languages are declarative, i.e., programming with these languages means describing the world. Using the programs means asking questions about the previously described world. Based on the answers to the questions from the logic reasoning engine, according to the world description, we can derive new facts. Figure 2 shows an example of how logic programming works.
Logic programming allows the programmer to specify the rules and facts, enabling the Prolog interpreter to infer answers to the given queries automatically. Here we explain some key concepts:
Program. A Prolog program consists of two parts: a list of facts () and a list of rules (). Throughout the paper, we use the over-tilde notation to denote a list of items. For example, refers to a list of entities, i.e., .
@@
An example is , which means horses are bigger than cats. Another example is , which means that the Gunzon Prize is in the list of prizes awarded to Haruki Murakami.
Rules. A Prolog rule is a Horn clause that comprises a head predicate and a list of body predicates placed on the left and right side of the arrow symbol ( :– ). A rule means that the left-hand side is logically implied by the right-hand side. The rule bodies are either positive or negative relations, corresponding to the requirements upon the presence or absence of facts. We use “” and “” as abbreviations for “” and “”, respectively. It is denoted as:
An example is , which means the relation is transitive. Another example is , which means is an inverse relation of . The last example here is , which is if there exists at least one that is a member of both and . It is a composite type of two predicates.
Queries. A query has the same structure as the body of a rule, i.e., it is a sequence of predicates separated by commas and executed against a database of facts. The logic reasoning engine will answer if the sequence of predicates in the query is according to the facts and rules. Otherwise, it will answer .
An example query is , which means asking the logic reasoning engine whether elephants are bigger than cats. Another example is , which means asking if the awards won by Haruki Murakami and Bob Dylan have overlapped.
Reasoning Rules. As shown in Figure 2, generating new facts through logic programming requires facts (Equation (missing) 2), rules (Equation (missing) 3), queries, and answers to the queries. To simplify the notation of this process, we bring up the concept of reasoning rules in this paper, which describes the inference process of using facts and rules (predicates) to reach the conclusion (a new fact in the form of a predicate) by omitting the process of querying and analyzing the query answers. A reasoning rule is denoted in this form:
@@ 3. Motivating Example
Figure 3 shows a motivating example of Drowzee. Assume we have the facts about whether Haruki Murakami and Bob Dylan have won the Nobel Prize, as illustrated in the left sub-figure. The question to ask LLMs is straightforward: We can ask whether Haruki Murakami/Bob Dylan has won the Nobel Prize or not. Asking and verifying this knowledge requires no logic reasoning. However, the straightforward questions are often not enough to unveil hallucinations. Therefore, more diversified questions (questions with intertwined and complex information, as illustrated in the right sub-figure) are needed.
In order to generate more diversified benchmarks, previous research (Yu et al., 2024; Li et al., 2023a) involves human experts to generate the questions and annotate the answers for hallucination checking. Although the manually generated benchmarks can unveil certain hallucinations, they suffer from several drawbacks. The landscape of knowledge is dynamic, with new information continuously surfacing and older information becoming obsolete. If facts change continuously over time, for instance Haruki Murakami were to win the Nobel Prize in the future, this would necessitate regular updates and corrections to the ground truth in existing datasets to reflect them. However, maintaining the accuracy of these benchmarks demands a significant amount of manual labor. Additionally, the quality of the questions might be inconsistent due to the differences in the experience and skills of the human experts who create them. Consequently, the efficiency and soundness of the manually generated benchmarks are not guaranteed.
@@
These two challenges can both be addressed by leveraging logic programming. We can derive new logically sound facts based on existing knowledge. With the new facts, we can then generate diverse questions and their ground truth answers. With the ground truth answers, we can generate test oracles to capture hallucinations. In short, the idea of using logic programming to tackle the challenges motivates the design of Drowzee.